A particle with a charge of -1.24×10^8 C is moving with instantaneous velocity v= (1.49×10^4 m/s)i + (-3.85×10^4 m/s)j. 

What is the force exerted on this particle by a magnetic field:

(A) B=(1.40 T)i + (-2.30 T)k and;

(B) B=(1.25 T)i + (2.00 T)j + (3.50 T)k?
Answer #1

magnetic force acting on the charge is given by F = 9 (TXB) where, 9 = -1.24 X108 v = (1.49x104) i + (-3.85x104) ģ(W/s) a) fo一个 -4 64 Ê IN : F-9 (TXB) = -1.24 x 108 (8.855X104 î + 3.427810* Î +5.39104 ķ F =-11.0X10î - 4.25X10ĵ- 6.68x10 N This is th

> Thankkkkk youuuuuu soooo mucchhhhhhh. Haha you saved my life

T1N4CH01 Tue, Jun 22, 2021 5:46 AM